# Zero-Shot Head Swapping in Real-World Scenarios

### v1

@@ 1 Introduction
Unlike face swapping, head swapping combines the entire head—including the face ID, face shape, and hairstyle—of the head image with the body image’s body, as their differences are highlighted in Figure 2. Therefore, while face swapping only applies the face ID to the body image, head swapping needs to seamlessly merge the entire head from the head image with the body in the body image. Due to its increased complexity, head swapping still faces numerous unresolved challenges. 1) For a realistic head swap, it requires seamless blending between the swapped head and the original body. 2) Unlike face swapping, where only the face ID features of the head image are extracted, head swapping have to also handle the structural information of the head that includes a hairstyle and a facial structure.
@@
These challenges stem from the use of face-centered cropped datasets with primarily frontal-facing views and masking methods. As shown in Figure 3(a), most head-swapping methods [33, 2, 14] perform head swapping only on face-centered cropped data, which may not fully contain the hair. Additionally, the masking methods proposed in [33, 2, 14] are optimized for these types of images and, thus lacking the ability to create a seamless transition between the head and body. Therefore, if users require head swapping in images that include the full head as well as the body, previous methods need an additional phase to paste the head back into the body, often resulting in inharmonious images. As this issue is clearly shown in Figure 4, if the original head in the body images has long hair beyond the cropped area, remnants of the hair remain visible. Furthermore, datasets with primarily frontal-facing views lack robustness to diverse facial orientations. Although there have been attempts to address this issue, such as HeSer, it operates in a few-shot manner  [41] by utilizing a wide range of view images extracted from video. This highlights the need for a more challenging dataset that includes upper body images with diverse facial orientations and has only a single image per person. To the best of our knowledge, no existing method efficiently addresses these real-world complexities in a zero-shot manner.
Therefore, we design a novel zero-shot head swapping approach, named Head Injection Diffusion (HID), with proposing Inverse-Orthogonal Mask (IOMask) and hair injection module, which are components of our approach to perform head swapping in real-world scenarios with this challenging dataset. Our approach not only handles diverse facial orientations in a zero-shot setting but also leverages a SHHQ [12] dataset that showcases the entire upper body. This approach combines the strengths of existing datasets while addressing the limitations of previous head swapping methods as illustrated in Figure 3. To tackle the remaining challenges in ensuring that the head of the head image blends naturally onto the body of the body image, we propose our IOMask, which automatically generates a context-aware mask. In head swapping tasks, automating the mask generation process is essential for real-world applications. Aligning the orientation of the head of the head image with head of the body image is crucial, but it is challenging for users to predict and provide such a mask manually and existing automatic masking methods have limitations as aforementioned. We address this issue the IOMask, extracted within HID, is obtained by leveraging the orthogonal component between the reconstruction condition and the editing condition, both of which originate from the inverted latent. Furthermore, to preserve face ID, face shape and hairstyle, we build our approach PhotoMaker V2 [1]  [^1] The code has been released but the paper has not been published yet., a state-of-the-art identity-oriented personalization model. Inspired by this method, we introduce our hair injection module to transfer the hair information more precisely.
@@ 2.1 Head Swap
Although there is relatively sparse research on head-swapping tasks, existing studies can be divided into two categories: few-shot approaches [41, 33] and zero-shot approaches [45, 14, 2]. Few-shot methods, such as those proposed by  [41, 33], utilize videos to give a few shots of head input from different views so that the head in the head image is better aligned with the head in the body image. On the other hand, HSDiffusion [45] is a zero-shot diffusion-based method but just aligns the center points of the head and body images and cutting and pasting the components accordingly. In other words, it does not align the direction of the head in the head image with the direction of the head in the body image. Therefore, HSDiffusion does not precisely align with traditional head swapping task. Other zero-shot approaches, like FaceX [14] and REFace [2], also leverage diffusion models for head swapping. FaceX is a unified method for handling diverse facial tasks and REFace is a face-swapping method adapted to support head swapping.
@@ 2.3 Image Editing
Figure 5: Overview of HID. Our HID consists of two main stages (left and right). In the left stage (blue region), we obtain updated text embeddings by fusing embeddings. Specifically, embeddings are respectively extracted from ID/Hair encoder and CLIP image encoder in each module and these embeddings are fused with corresponding text embeddings. These fused ID embeddings and fused hair embeddings replace the part of original text embeddings, resulting in updated text embeddings. In the right stage (white region), the final output $I_{o}$, a head swapped image, is generated. DDIM inversion is performed to reconstruct the image while leveraging our IOMask to infer which parts of the body image should be removed, thereby generating the head image. During this process, the updated text embeddings obtained in the left stage, along with the output from ControlNet, serves as conditioning inputs for the diffusion model.
@@ 3 Method
We propose a novel approach, HID, that takes a body image $I_{b}$ and a head image $I_{h}$ as inputs and outputs a swapped image $I_{o}$, where the head in $I_{b}$ is seamlessly replace with the one in $I_{h}$. We also introduce IOMask for precise mask extraction to guide head integration (Sec.3.2) and utilizes hair injection module to generate heads that match the input head images (Sec.3.3). Together, these components form our cohesive head-swap method (Sec.3.4). Figure 5 provides an overview of our process.
@@ 3.2 IOMask
For accurate head-swapping, obtaining a mask for head region is crucial to generate on the body image $I_{b}$. Inspired by [9, 28, 15], we achieve this by extracting the mask within our model. Given an body image $I_{b}$, we first apply DDIM [42] Inversion with our model, obtaining the latent $\hat{z}_{t}$ at a specific time step $t$. If we use DDIM Inversion with guidance scale $w=1$, that provides a faithful reconstruction when adapt denoising guidance scale the same. For specific time-step $t$, we can get $\epsilon_{\theta}(\hat{.}{z}_{t},t,C_{b})$ from denoising $\hat{z}_{t}$, where $C_{b}$ means body image’s text embeddings, and established this for a baseline. The Inverse-Orthognal map, called IO map, is then determined by the areas that differ under head-conditioned predicted noise ${\epsilon_{\theta}}(\hat{z}_{t},t,C_{h},\varnothing)$, where $C_{h}$ means head image’s text embeddings with fused ID embeddings and fused hair embeddings. Note that CFG is conducted for head-conditioned predicted noise. For simplicity, we’ll omit notation $\hat{z}_{t}$ and $t$.
@@ 4.2 Results
As shown in Table 1, our HID outperforms REFace [2] in both LPIPS and CLIP-I metrics for the head and hair regions. This result indicates that HID more accurately preserves ID and hairstyle in head images, which is essential for the head-swapping task. Moreover, we outperformed REFace on the FID score. This demonstrates that our generated images have superior quality, achieving a more natural integration of head and body features compared to existing methods.
@@
As shown in Figure 2, the swapped head must preserve key attributes of the head image, including skin tone, hairstyle, identity, and face shape. Figure 1 illustrates how our method effectively achieves these requirements, performing the head swap task while maintaining these essential characteristics. Additionally, thanks to the IOMask, we can generate appropriate masks that allow the creation of long hair where needed, ensuring that only the relevant head region is modified while leaving other areas unchanged. In Figure 7, we demonstrate that our method outperforms REFace [2], particularly in challenging cases where head orientation and gender differ significantly. Unlike REFace, which suffers from limitations inherent to the cropping approach—such as hair being cut off and not blending well with surrounding areas—our approach overcomes these issues, maintaining a seamless and coherent appearance even under these challenging conditions. Furthermore, unlike the mask used in REFace, our proposed IOMask effectively removes only the necessary face and hair areas from the body image, enabling a smooth and natural swap.
@@ 4.3.1 IOMask
We demonstrate the effectiveness of our IOMask through an ablation study, comparing the results of three configurations: the naive IO map, IO map without orthogonal components, and the full IO map. The naive IO map is defined as $\epsilon_{\theta}=\epsilon_{\theta}(C_{h},\varnothing)-\epsilon_{\theta}(C_{b},\varnothing)$, where the ID embeddings of both head and body conditions are without further refinement. The IO map without orthogonal components is defined as $\epsilon_{\theta}=\epsilon_{\theta}(C_{h},\varnothing)-\epsilon_{\theta}(C_{b})$ , omitting the orthogonal filtering step present in our full IO map. We visualized the IO map by overlaying it on body images, with red indicating regions of high values. As illustrated in Figure 8, the naive IO map introduces substantial randomness and often covers irrelevant areas. In contrast, both the IO map without orthogonal components and the full IO map provide more relevant coverage. Notably, the full IO map focuses more precisely on the regions that have to be swapped.
@@
Table 1: Quantitative comparison.
4.4 HID
We evaluate the performance impact of each element in the HID to show the effect of our methods. So, we sequentially remove the hair injection module and IO Mask, while applying both model with ControlNet and same input text, ’A man/woman img with hairstyle.’
@@ 5 Conclusion
We propose the HID approach, an efficient head swapping method aimed at real-world applications. Our approach allows the head image to blend seamlessly with the body image while preserving original details, resulting in a natural outcome. Despite these significant improvements over previous approaches, there are still areas that require refinement. For example, regions beyond the head, like the hands, neck, or other skin areas, may experience unintended changes. In the Figure 1, the second example from the right shows a case where the hands were masked to adjust their color, allowing them to match the skin tone of the head image. However, while the skin tone was transferred successfully, the pose and fine details of the original body image’s hands were not preserved, which highlights a limitation of our approach. Unlike previous methods that rely heavily on cropping or head-centric masking techniques, our approach takes into account which areas should be edited when conditioned with the head image, allowing for more flexibility and potential for improvement. Nevertheless, our method shows potential for further advancements, and future work could focus on better preserving details of these regions while still achieving seamless skin tone transfer.
@@ Acknowledgement
We appreciate the high-performance GPU computing support of HPC-AI Open Infrastructure via GIST SCENT.

### v2

@@ 1 Introduction
Unlike face swapping, head swapping combines the entire head, including the face ID, face shape, and hairstyle, of the head image with the body image’s body, as their differences are highlighted in Fig. 2. Therefore, while face swapping only applies the face ID to the body image, head swapping needs to seamlessly merge the entire head from the head image with the body in the body image. Due to its increased complexity, head swapping still faces numerous unresolved challenges. 1) For a realistic head swap, it requires seamless blending between the swapped head and the original body. 2) Unlike face swapping, where only the face ID features of the head image are extracted, head swapping have to also handle the structural information of the head that includes a hairstyle and a facial structure.
These challenges stem from the use of face-centered cropped datasets with primarily frontal-facing views and masking methods. As shown in  Fig. 3(a), most head-swapping methods [35, 2, 15] perform head swapping only on face-centered cropped data, which may not fully contain the hair. Additionally, the masking methods proposed in [35, 2, 15] are optimized for these types of images and, thus lacking the ability to create a seamless transition between the head and body. Therefore, if users require head swapping in images that include the full head as well as the body, previous methods need an additional phase to paste the head back into the body, often resulting in inharmonious images. As this issue is clearly shown in  Fig. 4, if the original head in the body images has long hair beyond the cropped area, remnants of the hair remain visible. Furthermore, datasets with primarily frontal-facing views lack robustness to diverse facial orientations. Although there have been attempts to address this issue, such as HeSer, it operates in a few-shot manner  [44] by utilizing a wide range of view images extracted from video. This highlights the need for a more challenging dataset that includes upper body images with diverse facial orientations and has only a single image per person. To the best of our knowledge, no existing method efficiently addresses these real-world complexities in a zero-shot manner.
@@
Therefore, we design a novel zero-shot head swapping approach, named Head Injection Diffusion (HID), with proposing Inverse-Orthogonal Mask (IOMask) and hair injection module, which are components of our approach to perform head swapping in real-world scenarios with this challenging dataset. Our approach not only handles diverse facial orientations in a zero-shot setting but also leverages a SHHQ [12] dataset that showcases the entire upper body. This approach combines the strengths of existing datasets while addressing the limitations of previous head swapping methods as illustrated in  Fig. 3. To tackle the remaining challenges in ensuring that the head of the head image blends naturally onto the body of the body image, we propose our IOMask, which automatically generates a context-aware mask. In head swapping tasks, automating the mask generation process is essential for real-world applications. Aligning the orientation of the head of the head image with head of the body image is crucial, but it is challenging for users to predict and provide such a mask manually and existing automatic masking methods have limitations as aforementioned. We address this issue by introducing the IOMask, extracted within HID, is obtained by leveraging the orthogonal component between the reconstruction condition and the editing condition, both of which originate from the inverted latent. Furthermore, to preserve face ID, face shape and hairstyle, we build our approach PhotoMaker V2 [1]  [^1] The code has been released but the paper has not been published yet., a state-of-the-art identity-oriented personalization model. Inspired by this method, we introduce our hair injection module to transfer the hair information more precisely.
@@ 2.1 Head Swap
Although there is relatively sparse research on head-swapping tasks, existing studies can be divided into two categories: few-shot approaches [44, 35] and zero-shot approaches [48, 15, 2]. Few-shot methods, such as those proposed by  [44, 35], utilize videos to give a few shots of head input from different views so that the head in the head image is better aligned with the head in the body image. On the other hand, HSDiffusion [48] is a zero-shot diffusion-based method but just aligns the center points of the head and body images and cutting and pasting the components accordingly. In other words, it does not align the direction of the head in the head image with the direction of the head in the body image. Therefore, HSDiffusion does not precisely align with traditional head swapping task. Other zero-shot approaches, like FaceX [15] and REFace [2], also leverage diffusion models for head swapping. [15] is a unified method for handling diverse facial tasks and [2] is a face-swapping method adapted to support head swapping.
@@ 2.3 Image Editing
Figure 5: Overview of HID. Our HID consists of two main stages (left and right). In the left stage (blue region), we obtain updated text embeddings by fusing embeddings. These fused ID embeddings and fused hair embeddings replace the part of original text embeddings, resulting in updated text embeddings. In the right stage (white region), the final output $I_{o}$, a head swapped image, is generated. DDIM inversion is performed to reconstruct the image while leveraging our IOMask to infer which parts of the body image should be removed, thereby generating the head image. During this process, the updated text embeddings obtained in the left stage, along with the output from ControlNet, serves as conditioning inputs for the diffusion model.
@@ 3 Method
We propose a novel approach, HID, that takes a body image $I_{b}$ and a head image $I_{h}$ as inputs and outputs a swapped image $I_{o}$, where the head in $I_{b}$ is seamlessly replaced with the one in $I_{h}$. We also introduce IOMask for precise mask extraction to guide head integration (Sec.3.2) and utilizes hair injection module to generate heads that match the input head images (Sec.3.3). Together, these components form our cohesive head-swap method (Sec.3.4).  Fig. 5 provides an overview of our process.
@@ 3.2 IOMask
For accurate head-swapping, obtaining a mask for head region is crucial to generate on the body image $I_{b}$. Inspired by [9, 30, 16], we achieve this by extracting the mask within our model. Given an body image $I_{b}$, we first apply DDIM [45] Inversion with our model, obtaining the latent $\hat{z}_{t}$ at a specific time step $t$. If we use DDIM Inversion with guidance scale $w=1$, that provides a faithful reconstruction when adapting the denoising guidance keep the scale same. For specific time-step $t$, we can get $\epsilon_{\theta}(\hat{.}{z}_{t},t,C_{b})$ from denoising $\hat{z}_{t}$, where $C_{b}$ means body image’s text embeddings, and established this for a baseline. The Inverse-Orthognal map, called IO map, is then determined by the areas that differ under head-conditioned predicted noise ${\epsilon_{\theta}}(\hat{z}_{t},t,C_{h},\varnothing)$, where $C_{h}$ means head image’s text embeddings with fused ID embeddings and fused hair embeddings. Note that CFG is conducted for head-conditioned predicted noise. For simplicity, we’ll omit notation $\hat{z}_{t}$ and $t$.
@@ 4.2 Results
As shown in  Tab. 2, our HID outperforms REFace [2] in both LPIPS and CLIP-I metrics for the head and hair regions. This result indicates that HID more accurately preserves ID and hairstyle in head images, which is essential for the head-swapping task. Moreover, we outperformed REFace on the FID score. This demonstrates that our generated images have superior quality, achieving a more natural integration of head and body features compared to existing methods.
@@
As shown in Fig. 2, the swapped head must preserve key attributes of the head image, including skin tone, hairstyle, identity, and face shape. Fig. 1 illustrates how our method effectively achieves these requirements, performing the head swap task while maintaining these essential characteristics. Additionally, thanks to the IOMask, we can generate appropriate masks that allow the creation of long hair where needed, ensuring that only the relevant head region is modified while leaving other areas unchanged. In Fig. 7, we demonstrate that our method outperforms REFace [2], particularly in challenging cases where head orientation and gender differ significantly. Unlike REFace, which suffers from limitations inherent to the cropping approach—such as hair being cut off and not blending well with surrounding areas—our approach overcomes these issues, maintaining a seamless and coherent appearance even under these challenging conditions. Furthermore, unlike the mask used in REFace, our proposed IOMask effectively removes only the necessary face and hair areas from the body image, enabling a smooth and natural swap.
@@ 4.3.1 IOMask
We demonstrate the effectiveness of our IOMask through an ablation study, comparing the results of three configurations: the naive IO map, IO map without orthogonal components, and the full IO map. The naive IO map is defined as $\epsilon_{\theta}=\epsilon_{\theta}(C_{h},\varnothing)-\epsilon_{\theta}(C_{b},\varnothing)$, where the ID embeddings of both head and body conditions are without further refinement. The IO map without orthogonal components is defined as $\epsilon_{\theta}=\epsilon_{\theta}(C_{h},\varnothing)-\epsilon_{\theta}(C_{b})$ , omitting the orthogonal filtering step present in our full IO map. We visualized the IO map by overlaying it on body images, with red indicating regions of high values. As illustrated in  Fig. 8, the naive IO map introduces substantial randomness and often covers irrelevant areas. In contrast, both the IO map without orthogonal components and the full IO map provide more relevant coverage. Notably, the full IO map focuses more precisely on the regions that have to be swapped.
4.3.2 HID
We evaluate the performance impact of each element in the HID to show the effect of our methods as shown in Fig. 9. So, we sequentially remove the hair injection module and IO Mask, while applying both model with ControlNet and same input text, ’A man/woman img with hairstyle.’
@@ 4.3.2 HID
Table 1: Quantitative comparison. Ours outperforms REFace [2] across all metrics.
@@ 5 Conclusion
We propose the HID approach, an efficient head swapping method aimed at real-world applications. Our approach allows the head image to blend seamlessly with the body image while preserving original details, resulting in a natural outcome. Despite these significant improvements over previous approaches, there are still areas that require refinement. For example, regions beyond the head, like the hands, neck, or other skin areas, may experience unintended changes. In the Fig. 1, the second example from the right shows a case where the hands were masked to adjust their color, allowing them to match the skin tone of the head image. However, while the skin tone was transferred successfully, the pose and fine details of the original body image’s hands were not preserved, which highlights a limitation of our approach. Unlike previous methods that rely heavily on cropping or head-centric masking techniques, our approach takes into account which areas should be edited when conditioned with the head image, allowing for more flexibility and potential for improvement. Nevertheless, our method shows potential for further advancements, and future work could focus on better preserving details of these regions while still achieving seamless skin tone transfer.
@@ Acknowledgement
This work was supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)). This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2025-00555621). We appreciate the high-performance GPU computing support of HPC-AI Open Infrastructure via GIST SCENT.
@@
Figure 10: Additional qualitative comparisons between InstantID and Ours, as well as Ours with Sapiens mask and Ours.
@@ 6.2 IOMask
7 Problem Definition
Face swapping can be defined in various ways, such as replacing only the eyes, nose, and mouth while preserving the original face shape. In our case, we define a “head swap” as the task of swapping only the head, which is considered part of the human body. Consequently, long beards should be reflected in the swap, whereas accessories should not. To ensure that accessories are not swapped, we refined the training dataset accordingly. As shown in Fig.7 in manuscript, facial hair is partially preserved in the swap to a reasonable extent. Moreover, for a seamless transition, it is necessary to process not only skin of the neck area, which connects to the head, but also the hands. Our IOMask has the potential to enable this process as shown in Fig. 10, first row.
Head (crop) Hair (crop) ID sim $\uparrow$ Head Pose $\downarrow$ Recon. $\uparrow$
[…]
REFace 0.4932 0.7430 0.3189 0.8312 0.5197 20.33 0.6397
InstantID* 0.4802 0.8090 0.3251 0.8470 0.2821 20.64 0.6308
Ours 0.4567 0.8615 0.2966 0.8644 0.5546 21.66 0.6286
Table 2: Quantitative comparison.
8 Analysis on Qualitative Results
The reason parts other than the head are sometimes altered, called hallucinations or artifacts, is that the mask leaves those areas uncovered. As mentioned in the Sec.1, if we want to generate a head independent of the original body’s head, using predicted noise’s difference methods are effective. However, these methods have a drawback: they often open up areas that do not need to be changed. IOMask helps minimize these issues while still preserving the advantages of such approaches as shown in Fig. 10, second and third row. Note that REFace is a method that uses the crop-and-paste-back approach, ensuring that other parts of the image remain unchanged which has its own limitations discussed in Sec.1. When using an off-the-shelf mask, it only fills the head region of $I_{b}$, which can lead to issues such as the inability to generate long hair or the appearance of sharp skin boundaries, as shown in Fig.10. In contrast, IOMask allows for more flexible and natural generation.
We use a dataset that includes the entire upper body. Therefore, rather than providing only head orientation information, we utilize OpenPose, which offers richer information to ensure better alignment with the body. It is also worth noting that, as shown in Fig.5, openpose has difficulty capturing fine details of hands [13]. Since our focus is on head swapping, this limitation lies outside the scope of our work. However, it is an important aspect to consider for future improvements, particularly in full-body swap scenarios.
9 Additional Quantitative Comparisons
We additionally evaluated our method using four metrics: ID similarity with the face recognition model [10] head pose L2 distance with the head pose estimation network, HopeNet [41], LIPIS and CLIP-I for cropped images to compare with REFace, as well as a reconstruction metric, MSE, to assess the degree of preservation outside of the head area. Furthermore, we conducted an additional comparison with SDXL Inpainting InstantID [49], using Sapiens [23] mask, with openpose ControlNet. InstantID not only struggles to preserve identity compared to our method but also fails to transfer hair. Our method outperformed in all aspects except for head pose and reconstruction. The weaknesses observed in head pose and reconstruction can be attributed to the trade-off between flexibility and preservation, which arises from the different masking methods.